OpenStudy (anonymous):

Somebody PLEASE help? Determine the factors of x2 − 8x − 12.

OpenStudy (ganpat):

@jiteshmeghwal9 : -6 * -2 should be equal to + 12.. right ??

OpenStudy (jiteshmeghwal9):

oops sorry!!!!!!!!

OpenStudy (ganpat):

@Rammsey: is the question correct ?? with signs ??

OpenStudy (jiteshmeghwal9):

yeah! i also think that either question has some mistake or the solution is no factors:/

OpenStudy (ganpat):

factor = -b +- sq rt (b2 - 4ac)... @jiteshmeghwal9 : You remember some formula like this ??

OpenStudy (jiteshmeghwal9):

ya!

OpenStudy (ganpat):

\[factor = -b +- \sqrt{(b ^{2} - 4ac)}...\]

OpenStudy (jiteshmeghwal9):

\[\huge{\frac{-b \pm \sqrt{b^2-4ac}}{2a}}\]

OpenStudy (jiteshmeghwal9):

u can find alpha & beta roots of the equation by quadratic formula & then the factors would be in the form as\[(x-\alpha)(x-\beta)=0.\]

OpenStudy (jiteshmeghwal9):

so just solve & u can reach ur answer 100%.

OpenStudy (ganpat):

comparing it with ax2 + bx + c x2 − 8x − 12. a = 1, b = -8 , c = -12 So, \[8 +- \sqrt{(64 - 4(1) (-12))} / 2(1) = 8 +- \sqrt{(112)} = 8 +- 10.58\] so the factors are: (8 + 10.58) / 2 and (8 - 10.58)/2 18.58/2 and -2.58/2 so the factors are , 9.29 and -1.29...

OpenStudy (jiteshmeghwal9):

yeah so the factors are (x-9.29) & (x+1.29) @Ganpat & @Rammsey :D

OpenStudy (ganpat):

cool !!! thanks :) @jiteshmeghwal9

OpenStudy (jiteshmeghwal9):

yw:)